For a platinum wire, the coefficients of variation of resistance with temperature are α=3.94×10^(–3) °C^(–1)
and β=–5.82×10^(–7) °C^(–2).
A thermometer is fabricated using
this wire with
R0=12.00 Ω
where
R0 signifies the resistance of the material of wire
at ice point. When the thermometer is placed in contact with a heat bath, the
resistance is found to be
15.20Ω.
Calculate the temperature of the bath.
1
Expert's answer
2020-02-25T10:20:47-0500

Solution. The temperature dependence of resistance on temperature can be represented by the formula


"R_t=R_0(1+\alpha t+\beta t^2)"

where Rt is resistance at temperature t0C ; R0 is resistance at temperature 00C;  α=3.94×10^(–3) °C^(–1) and β=–5.82×10^(–7) °C^(–2) are constants. (the formula is valid in the temperature range 00С-8500С) According to the conditions of the problem Rt=15.20Ω,

R0=12.00 Ω. As result get quadratic equation


"15.20=12.00(1+3.94\u00d710^{\u20133} t - 5.82\u00d710^{\u20137} t^2)""- 5.82\u00d710^{\u20137} t^2+3.94\u00d710^{\u20133} t - 0.267=0"


Find the roots of the quadratic equation


"D=(3.94×10^{–3})^2-4( - 5.82×10^{–7} )(-0.267) \approx (3.86×10^{–3})^2"

"t_1=\frac {-3.94\times10^{-3}-3.86\times10^{-3}}{-2\times 5.82\times 10^{-7}}\approx 6692^0C"

The temperature t1 does not belong to the interval 00С-8500С.


"t_2=\frac {-3.94\times10^{-3}+3.86\times10^{-3}}{-2\times 5.82\times 10^{-7}}\approx 69^0C"

Answer. 690C